The base of a triangle is one-half as long as its height. If the area of the triangle 100 yards, find its height
Answer by TimothyLamb(4379)

a = hb/2
b = h/2
100 = hb/2
100 = hh/(2*2)
hh = 400
h = sqrt( 400 )
h = 20 yards
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---